OpenStudy (cyclone3260):

Select all of the following points that lie on the graph of f(x) = 7 - 3x. (-2, 1) (-1, 10) (0, 7) (1, 5) (2, 1)

OpenStudy (anonymous):

(-1, 10) f(x) = 7 - 3x. f(-1) = 7 - 3(-1) = 10.

OpenStudy (cyclone3260):

Any more? There's supposed to be more than 1.